# On Unbiased Low-Rank Approximation with Minimum Distortion

### v1

@@ 1 Introduction
The problem of unbiased low-rank approximation has recently been considered in the machine learning literature, and in particular the authors [2] show that a duality-based lower bound on the expected distortion can be matched by a complicated sampling procedure. Their procedure is only explicitly defined for the co-rank one scenario $r+1=\mathsf{rank}(P)$. For arbitrary $r$, they point out that an algorithm can be written down by recursively applying the co-rank one version (see the section titled Proof of Lemma A.6 in the supplementary material of [2]). Our main contribution in the present note is to show that a simple sparsification procedure applied to the singular components of $P$ achieves this lower bound for any $r$, and therefore gives a simpler and more direct sampling procedure for producing optimal, unbiased, low-rank approximations.
@@
each with equal probability 1/2. Both $Q,Q^{\prime}$ achieve the minimal distortion $\mathbb{E}\|P-Q\|_{F}^{2}=\mathbb{E}\|P-Q^{\prime}\|_{F}^{2}=8$, as does the mixture $Q_{t}$ which takes the values $Q_{t}=Q$ with probability $t$ and $Q_{t}=Q^{\prime}$ with probability $1-t$ for any $0\leq t\leq 1$. This variety of optimal solutions is in stark contrast to the rigid description of optimizers for the vector valued case [1], and suggests there is a rich space of possible optimal unbiased approximations to explore.
@@ 2 The Sampling Procedure
Of the remaining $N-k$ components, sample a subset $I$ of $r-k$ elements of the indices $\{k+1,\ldots,N\}$ with inclusion probabilities proportional to their $d_{i}$ values. There are many ways to do this sampling, and we describe one simple way in the following paragraph (see also [1, 10]). Let’s then set the resulting $Q^{\prime}=\mathsf{diag}(Q^{\prime}_{1},\ldots,Q^{\prime}_{N},0,\ldots,0)$ by the components
@@
As promised, one way to sample the index set $I$ is as follows. Divide the interval $[0,r-k]$ up into $N-k$ disjoint segments, with segment $i$ (for $i=k+1,\ldots,N$) having length proportional to $d_{i}$. Then randomly generate a single number $S\sim\mathsf{unif}[0,1]$, and if $S+j$ is in segment $i$ for some $j=0,1,\ldots,r-k-1$, then include the index $i$ in $I$. Otherwise, do not include that index $i$ in $I$. It may be desirable to randomly permute the order the segments, but this is not required to achieve the correct inclusion probabilties.
One can verify this gives the desired inclusion probabilities by zooming in on a single, unit-length subinterval $[z,z+1],z\in\mathbb{Z}$ of $[0,r-k]$. If this subinterval entirely contains segment $i$, then segment $i$ will be included with probability equal to its length (which is proportional do $d_{i}$). If part of segment $i$ lies in an adjacent subinterval, then the sum of the lengths of the two parts will be the inclusion probability (this uses the fact that no segment can have length more than one, so the event that both parts are chosen cannot happen). No segment can intersect more than two subintervals with nonzero measure, since their lengths are bounded by one.
@@
9:$S\leftarrow\mathsf{unif}[0,1]$
10:$I\leftarrow\{\}$
11:$\mathsf{length}\leftarrow 0$
12:$C\leftarrow\frac{r-k}{\sum_{i=k+1}^{N}d_{i}}$
@@
20:$Q^{\prime}=\mathsf{diag}(Q^{\prime}_{1},\ldots,Q^{\prime}_{N},0,\ldots,0)$ where $Q_{i}^{\prime}=\begin{cases}d_{i}&\;,\;i=1,\ldots,k\\ c=\frac{1}{C}=\frac{\sum_{j=k+1}^{N}d_{j}}{r-k}&\;,\;i\in I\\ 0&\;,\;\mathsf{otherwise}\end{cases}$

### v2

@@ 1 Introduction
The problem of unbiased low-rank approximation has recently been considered in the machine learning literature, and in particular the authors [2] show that a duality-based lower bound on the expected distortion can be matched by a complicated sampling procedure. In [2], their procedure is explicitly defined for the co-rank one scenario $r+1=\mathsf{rank}(P)$, and for arbitrary $r$ they point out that an algorithm can be written down inductively in $r$ (see the section titled Proof of Lemma A.6 in the supplementary material). Our main contribution in the present note is to show that a simple sparsification procedure applied to the singular components of $P$ achieves this lower bound for any $r$, and therefore gives a simpler and more direct sampling procedure for producing optimal, unbiased, low-rank approximations.
@@
each with equal probability 1/2. Both $Q,Q^{\prime}$ achieve the minimal distortion $\mathbb{E}\|P-Q\|_{F}^{2}=\mathbb{E}\|P-Q^{\prime}\|_{F}^{2}=8$, as does the mixture $Q_{t}$ which takes the values $Q_{t}=Q$ with probability $t$ and $Q_{t}=Q^{\prime}$ with probability 1t1-t for any $0\leq t\leq 1$. This variety of optimal solutions is in stark contrast to the rigid description of optimizers for the vector valued case [1], and suggests there is a rich space of possible optimal unbiased approximations to explore.
@@ 2 The Sampling Procedure
Of the remaining NkN-k components, sample a subset $I$ of rkr-k elements of the indices $\{k+1,\ldots,N\}$ with inclusion probabilities proportional to their $d_{i}$ values. There are many ways to do this sampling, and we describe one simple way in the following paragraph (see also [1, 8]). Let’s then set the resulting $Q^{\prime}=\mathsf{diag}(Q^{\prime}_{1},\ldots,Q^{\prime}_{N},0,\ldots,0)$ by the components
@@
As promised, one way to sample the index set $I$ is as follows. Divide the interval [0,rk][0,r-k] up into NkN-k disjoint segments, with segment $i$ (for $i=k+1,\ldots,N$) having length proportional to $d_{i}$. Then randomly generate a single number $S\sim\mathsf{unif}[0,1]$, and if $S+j$ is in segment $i$ for some $j=0,1,\ldots,r-k-1$, then include the index $i$ in $I$. Otherwise, do not include that index $i$ in $I$. It may be desirable to randomly permute the order the segments, but this is not required to achieve the correct inclusion probabilties.
One can verify this gives the desired inclusion probabilities by zooming in on a single, unit-length subinterval $z,z+1],z\in\mathbb{Z}$ of [0,rk][0,r-k]. If this subinterval entirely contains segment $i$, then segment $i$ will be included with probability equal to its length (which is proportional do $d_{i}$). If part of segment $i$ lies in an adjacent subinterval, then the sum of the lengths of the two parts will be the inclusion probability (this uses the fact that no segment can have length more than one, so the event that both parts are chosen cannot happen). No segment can intersect more than two subintervals with nonzero measure, since their lengths are bounded by one.
@@
[…]
@@
[…]